OpenStudy (anonymous):

i am a 5th grader i need help can some body explain how long will it take me to travel 20 miles at the rate of 45mph

OpenStudy (anonymous):

idk but i am ur fan

OpenStudy (anonymous):

o.......k

OpenStudy (radar):

I may not be smarter than a 5th grader but here is my take on. Distance = Rate times Time. d=rt

OpenStudy (radar):

Now in your problem what is given? Rate is 45 mph. Distance is 20 miles. plug them in and solve.

OpenStudy (radar):

20=45mphxT, T=20milesd/45mph answer will be in hour unit

OpenStudy (radar):

20/45=4/9hr or 4/9*60=26.67min